# Improving governance outcomes through AI documentation: Bridging theory and practice

### v1

@@ 4.2. Potential Impacts & Challenges
4.2.1. Potential impact #1: Documentation could inform downstream stakeholders about effective and responsible use.
The most straightforward potential impact of documentation is that it can provide essential context, such as the original motivation for developing a dataset, model, or system, their intended use cases, and information on system properties necessary for effective and responsible deployment.
@@
Hypotheses. Documentation could inform downstream stakeholders by clearly articulating the intended uses of data, models, and systems and provide guidance on responsible deployment. Practitioners typically develop systems with specific use cases in mind, and other development teams need to understand these intended use cases to assess alignment with their goals. If aligned, teams can then modify the systems to meet their current objectives. Practitioners argue documentation can also clarify underlying assumptions, such as data collection methods and representativeness, which are critical for assessing the suitability of a dataset for specific purposes (Reid and Williams, 2023; Rostamzadeh et al., 2022; Zheng et al., 2022). Also, it can help identify potential deployment issues, such as mismatches between the original training data and the intended deployment context (Boyd, 2021).
@@
Moreover, documentation could assist downstream practitioners in making implementation decisions that support effective and responsible product development. For example, documentation could help practitioners find and compare system components aligned with their goals, including alternatives to AI approaches (Heger et al., 2022; Reid and Williams, 2023; Shen et al., 2022; Zheng et al., 2022). It might also provide statistical summaries of data (Holland et al., 2018) or or model performance evaluations such as area under the curve (AUC) (Mitchell et al., 2019) to help inform developers’ choices of how to train models or implement appropriate guardrails.
@@
Challenges. While practitioners acknowledge that the benefits of documentation could outweigh the costs of creating and maintaining it (Heger et al., 2022), when organizations fail to provide clear incentives for high-quality documentation, practitioners are less likely to prioritize it. This lack of motivation can hinder the realization of documentation’s potential benefits, but particularly its ability to inform downstream users.
One major challenge is the lack of organizational incentives for documentation. Organizations typically do not prioritize documentation unless regulations require it for compliance or clients specifically request it (Chang and Custis, 2022; Miceli et al., 2021; Heger et al., 2022). Practitioners often view producing high-quality documentation as less relevant to their evaluations and promotions than development tasks that directly contribute to product outputs (Chang and Custis, 2022; Heger et al., 2022). This disconnect between the benefits of documentation and practitioner objectives is further exacerbated when the practitioners who benefit most from documentation (i.e., consumers) are not the same as those responsible for producing it (i.e., producers). In organizations where governance processes mandate documentation, a lack of familiarity with AI documentation concepts among developers could potentially magnify tensions between product development and governance teams (Ali et al., 2023; Ahlawat et al., 2024). Moreover, practitioners are often more motivated by compliance than by normative considerations, which could lead to blind spots when novel ethical issues arise outside existing policies (Heger et al., 2022).
Facing time constraints and competing demands, practitioners may deprioritize documentation, sometimes cutting corners to save time (Chang and Custis, 2022; Heger et al., 2022). This can diminish the value of documentation artifacts, as evidenced by studies showing that practitioners sometimes reuse content from existing documentation, even when it pertains to different systems, leading to inadequate or incorrect documentation of their current system (Pushkarna et al., 2022; Heger et al., 2022). Practitioners also sometimes leave questions unanswered rather than seek out the necessary information (Heger et al., 2022) or prioritize recording information that is most valuable to them, potentially omitting critical details for others less familiar with the system (Hind et al., 2019). Furthermore, because practitioners often forget relevant information if they do not record it during development, their reconstruction of key system information after the fact can be time-consuming and error prone (Reid and Williams, 2023).
Studies of publicly available documentation artifacts frequently highlight issues with incomplete or incorrect information, which can mislead downstream stakeholders regarding the characteristics of AI systems or components (Pepe et al., 2024; Liang et al., 2024; Geiger et al., 2020). For example, many publicly available model cards lack details about out-of-scope uses, limitations, or environmental impacts (Pepe et al., 2024; Liang et al., 2024). Less than 30% of model cards contain evaluation results, a critical piece of information for downstream users (Liang et al., 2024). Also, some publicly available documentation contains incorrect license information, potentially leading users to inadvertently violate license terms (Pepe et al., 2024).
@@
4.2.2. Potential impact #2: Documentation could support cross-functional collaboration and communication about AI risks.
@@
Documentation has the potential to act as a bridge between different stakeholder groups involved in the AI development lifecycle, including machine learning engineers, data scientists, product managers, user experience designers, researchers, and legal teams. Since no single group oversees all stages of development, effective documentation can facilitate both indirect and direct communication among these diverse teams. Documentation artifacts can convey essential information across organizational boundaries, helping stakeholders who may not interact regularly to align their understanding of the system. Documentation can also prompt and support conversations among stakeholders directly, enabling them to work towards common goals despite their diverse backgrounds, frames of reference, or areas of expertise. This bridging function of documentation could break down organizational silos(Chmielinski et al., 2024) and promote effective collaboration across different teams (Crisan et al., 2022; Gilbert et al., 2023; Heger et al., 2022; Mohammad, 2022; Pushkarna et al., 2022; Raji and Yang, 2020; Richards et al., 2020; Shen et al., 2022; Srinivasan et al., 2021).
@@
Hypotheses. AI documentation could serve as an initial point of reference for downstream practitioners, offering a basic overview of the system that allows them to build foundational knowledge before engaging with developers for more in-depth discussions. Rather than acting as a comprehensive, self-contained repository of information, documentation artifacts could provide an initial layer of understanding, enabling downstream practitioners to identify relevant areas for further exploration and formulate more informed questions when collaborating with developers. This could foster more productive and targeted interactions between teams, enhancing their ability to work with or govern AI systems efficiently. Empirical research suggests that practitioners often approach documentation this way. For instance, UX professionals often prefer direct discussions with data scientists rather than relying solely on documentation artifacts (Liao et al., 2023). Similarly, many AI practitioners review existing documentation only briefly before seeking further clarification through meetings or discussions (Heger et al., 2022).
@@
In more robust forms, documentation can facilitate deeper deliberation among stakeholders about the system’s design and the documentation process itself. For example, documentation has been shown to aid teams in collectively reflecting on socially constructed concepts, such as gender, and making informed decisions about how to annotate such attributes in datasets (Pushkarna et al., 2022). Also, the process of creating documentation can help stakeholders identify and discuss trade-offs between competing objectives in system design (Shen et al., 2022).
@@
Challenges. Several challenges complicate the role of documentation in facilitating cross-functional collaboration. The modern AI supply chain, characterized by the development of general-purpose models that can be adapted to a variety of applications downstream (Vaswani et al., 2017), often limits opportunities for direct communication between stakeholders. In some cases, different teams within the same organization may adapt a general-purpose model for various specific tasks. In these instances, documentation might prompt the downstream team to directly reach out to the original development team to discuss the system’s capabilities and limitations. In other cases, a team deploying a model may rely on a model developed by a different organization (Jones, 2023). In this case, it will typically be more difficult if not impossible for the downstream deployment team to directly communicate with the original development team.
In either situation, the original developers may not fully anticipate the variety of downstream uses, leading to documentation that lacks critical details needed to identify potential risks or harms. Particularly when upstream and downstream teams cannot communicate directly, it will be more challenging for them to work together to understand the model’s risks within a given deployment context. This challenge is especially pronounced for less technical downstream stakeholders, who may struggle to understand the functionality and risks associated with these models.
More interactive forms of documentation, such as those that allow users to engage with the models directly or customize the level of detail presented, could help mitigate these issues by making the information more accessible and relevant (Crisan et al., 2022; Liao et al., 2023; Moore et al., 2023), especially when direct communication between teams is limited. However, further research is needed to determine the most effective ways to document general-purpose systems while avoiding overwhelming documentation consumers with excessive or irrelevant details,
Another major challenge is that documentation that is misaligned with the needs of its intended consumers can complicate rather than facilitate stakeholder communication. If documentation artifacts are overly technical or jargony, documentation consumers who are less technical or less familiar with the system may struggle to understand what details are most relevant, or misinterpret described practices (Crisan et al., 2022; Liao et al., 2023; Miceli et al., 2021; Shen et al., 2022); if documentation lacks relevant details, documentation consumers may overlook risks that become apparent only later in the development process, which may require teams to substantially backtrack. In either case, mismatch in comprehension can contribute to needless disruption to normal workflows. If practitioners perceive that documentation leads to seemingly egregious friction due to misunderstandings, they may be more prone to oversimplify the documentation they produce or gloss over relevant details that they worry will lead to confusion.
@@
4.2.3. Potential impact #3: Documentation could prompt practitioners to deliberate and act on the ethical impacts of AI
Researchers have suggested that documentation has the potential to encourage practitioners to consider the ethical implications of their systems, leading to more responsible development and use of data, models, and systems. By integrating ethical reflection into the documentation process, practitioners may become more aware of the potential harms their systems could cause, prompting them to make more informed decisions.
@@
Hypotheses. One hypothesis is that when downstream documentation consumers consult documentation on data, models, and systems, it can prompt consideration of potential harms. This consideration could lead practitioners to make more careful decisions about if and when to use certain systems or components. Yet, empirical support for this hypothesis is inconsistent. One study of 23 practitioners examined whether reviewing data documentation artifacts could encourage practitioners to notice ethical issues, contextualize them, and formulate risk mitigation plans (Boyd, 2021). One subset received a scenario description, dataset, and data documentation, while the rest received only the scenario description and dataset. The researcher found that those with data documentation were more likely to notice ethical issues unprompted, though those without it often recognized concerns when the researcher pressed them — but practitioners had difficulty formulating action plans regardless of whether they had access to documentation. These findings suggest that while documentation artifacts can aid ethical deliberation, documentation is not always necessary and likely not sufficient to enable mitigation of ethical issues.
@@
Documentation also might sensitize producers to the ethical impacts of their work. Documentation frameworks often ask documentation producers to consider and describe the ethical dimensions of their systems. For example, the one framework prompts practitioners to think about privacy implications by inquiring about sensitive information in data (Gebru et al., 2021). Another asks about known biases, ethical issues, and safety concerns (Arnold et al., 2019), and several frameworks encourage reflection on how systems might negatively impact marginalized users or populations (Brajovic et al., 2023; McMillan-Major et al., 2021; Shen et al., 2022).
@@
Challenges. The assumption that documentation artifacts and processes can promote ethical reasoning among practitioners hinges on the belief that practitioners are sufficiently aware of AI’s ethical risks to users, non-users, and society that they can recognize them when prompted. However, without proper training in responsible AI practices and exposure to groups potentially harmed by AI systems, practitioners may not connect documentation with ethical risks (Heger et al., 2022). Consequently, their consideration of harms may only partially encompass the scope of risks, especially those requiring non-technical mitigations. For example, practitioners often misunderstand how bias can manifest in their work, leading them to incorrectly state in their documentation that bias concerns are not relevant (Hind et al., 2019). While it may not be necessary for ethicists to be embedded in the work process for practitioners to make responsible choices, as some have suggested (McLennan et al., 2022), findings point to the need for more support and training for practitioners in how to identify, document, and effectively communicate potential risks (Heger et al., 2022; Madaio et al., 2024).
Furthermore, practitioners and organizations may resist documenting ethical impacts they have identifed. Publicly available documentation artifacts, such as Hugging Face model cards and GitHub repository README files, often show little consideration of ethical concerns (Pepe et al., 2024; Liang et al., 2024). While this omission might occur because practitioners are unaware of potential harms (Heger et al., 2022), their absence could also result from practitioners being hesitant to record ethical concerns they did recognize. Sometimes practitioners resist addressing ethical impacts because they feel unqualified to speculate on numerous potential use cases and their impacts (Crisan et al., 2022; Heger et al., 2022; Hind et al., 2019; Nunes et al., 2022). Others worry that detailing ethical concerns might give downstream stakeholders a false sense of security (Crisan et al., 2022). This hesitation could also stem from a desire to avoid personal responsibility for their actions (Nunes et al., 2022). Alternatively, practitioners could plausibly have concerns that documenting potential ethical harms could create a legal or public relations risk if outside stakeholders gained access to documentation.
Even if documentation effectively engages practitioners in ethical reflection, they may have little influence over their organization’s development goals, deliverables, and timelines. As a result, they might be unable to make the changes to data, models, or systems that they have identified as useful or necessary to address ethical risks through the documentation process. One ethnographic study found that business demands, not the beliefs of data subjects or practitioners, largely determined the organization’s documentation approaches (Miceli et al., 2021). For example, although practitioners who reflected on data labeling practices recognized that socially constructed identities are complex, the organization nevertheless chose to represent identity in a reductive way, such as by defining race according to discrete, mutually exclusive categories. The authors concluded that explicit and implicit power structures among internal organizational stakeholders significantly affect practitioners’ documentation practices and ability to shape outcomes. Documentation approaches must be responsive to these constraints. Otherwise, documentation about ethical considerations may not promote meaningful action or may be incomplete or misleading.
On the whole, evidence on the extent to which AI documentation frameworks can truly enhance ethical decision-making is mixed. Practitioners are capable of ethical deliberation about AI, but do not always thoroughly explore ethical issues or record the results in documentation artifacts. Even if practitioners engage in ethical deliberation when creating and using documentation, organizational environments can enable or constrain individual practitioners’ normative choices.
@@
4.2.4. Potential impact #4: Documentation could catalyze overall improvements in governance and development.
Researchers have suggested that documentation can significantly influence governance and development practices broadly. In practical settings, AI documentation doesn’t function independently; documentation processes typically require practitioners to engage in other activities that can improve AI outcomes. Thus, the documentation process can catalyze behaviors that enhance the quality of AI system development and governance beyond documentation processes and artifacts.
@@
Hypotheses. Framework authors posit that documentation could improve the rigor with which practitioners develop AI systems (Holland et al., 2018). By requiring practitioners to justify their development choices, documentation may lead to more careful decision-making (Bender and Friedman, 2018; Chmielinski et al., 2024; Holland et al., 2018). Also, clear and comprehensive documentation supports reproducibility, aiding other practitioners in retraining and validating systems consistently (Adkins et al., 2022; Baracaldo et al., 2022; Hind et al., 2019). Since AI practitioners are often driven by a commitment to scientific rigor (Winecoff and Watkins, 2022), documentation that emphasizes this aspect can serve as a ”value lever (Shilton, 2013),” further encouraging engagement with the documentation process .
@@
Documentation could also facilitate proactive risk mitigation, which is often more efficient than reactive approaches (Mitchell et al., 2019). When issues arise in already-deployed systems, organizations typically need to fix the problem without disrupting service quality or availability. They might need to roll back to a previous version of the system that is less performant or apply quick patches that may not fully resolve the issue. Documentation that helps identify potential problems before deployment can lead to more comprehensive solutions than addressing issues post hoc once they are discovered in production systems (Holland et al., 2018). And helping teams discover and use approved AI components that have already undergone relevant risk management processes rather than creating new ones can help ensure risks are not re-introduced and overlooked.
@@
Furthermore, documentation could preserve institutional knowledge. In any organization, particularly complex ones, consistently institutionalizing values and best practices can be challenging. Documentation could serve as a means of conveying technical information about systems, and the organization’s policies and values that influence development and use. For instance, documentation can aid in onboarding new employees by providing information about the systems they’ll work on and communicating the organization’s approach to development and governance (Adkins et al., 2022; Chang and Custis, 2022; Heger et al., 2022).
@@
Lastly, documentation could facilitate both internal and external audits (Brajovic et al., 2023; Crisan et al., 2022; Hupont and Gomez, 2022; Miceli et al., 2021). Audits may focus on the components of AI systems, the overall application, or the processes used in development (Mökander et al., 2023). Documentation could help organizations demonstrate that their claims—such as providing equitable system performance across different demographic groups—are backed by evidence. In these cases, documentation could guide auditors in selecting appropriate methods based on the auditor’s objectives and the available data. When audits examine the robustness of an organization’s governance, documentation becomes even more critical (Clavell, 2024). Without comprehensive documentation, organizations may struggle to demonstrate the integrity and consistency of their processes.

### v2

@@ 4.2. Theories of change
4.2.1. Theory of change #1: Documentation could inform downstream stakeholders about effective and responsible use.
@@
Documentation can provide essential background context on the system or component’s development, informing appropriate integration of data, models, methods, or systems into technologies. For example, practitioners typically develop systems with specific use cases in mind, and other development teams need to understand these intended use cases to assess alignment with their goals. If aligned, teams can then modify the systems to meet their current objectives. Practitioners argue documentation can also clarify underlying assumptions, such as data collection methods and representativeness, which are critical for assessing the suitability of a dataset for specific purposes (Reid and Williams, 2023; Rostamzadeh et al., 2022; Zheng et al., 2022). Also, documentation on background context can help identify potential issues prior to deployment, such as mismatches between the original training data and downstream practitioners’ intended deployment setting (Boyd, 2021).
@@
Moreover, documentation could provide information that helps downstream practitioners make sound implementation decisions. For example, documentation could help practitioners find and compare system components aligned with their goals, including alternatives to AI approaches (Heger et al., 2022; Reid and Williams, 2023; Shen et al., 2022; Zheng et al., 2022). It might also provide statistical summaries of data (Holland et al., 2018) or or model performance evaluations such as area under the curve (AUC) (Mitchell et al., 2019) to help inform developers’ choices of how to train models or implement appropriate guardrails.
@@
The type of information that is most useful to downstream stakeholders may depend on whether the documentation pertains to data, models, systems, or methods and tasks. For example, system documentation can help downstream stakeholders understand functionality and risks that emerge from the interactions of underlying components (Arnold et al., 2019; Procope et al., 2022). Data documentation, on the other hand, can help downstream practitioners assess the downstream impact of upstream inputs, such as potentially privacy-compromising data inputs or sampling methods that could result in unfair user treatment (Gebru et al., 2021; Boyd, 2021).
@@
4.2.2. Theory of change #2: Documentation could support cross-functional collaboration and communication about AI risks.
@@
Since no single group oversees all stages of development, effective documentation can facilitate both indirect and direct communication among these diverse teams. Documentation artifacts can convey essential information across organizational boundaries, helping stakeholders who may not interact regularly to align their understanding of the system through documentation artifacts. Documentation can also prompt and support conversations among stakeholders directly, enabling them to work towards common goals despite their diverse backgrounds, frames of reference, or areas of expertise. This bridging function of documentation could break down organizational silos (Chmielinski et al., 2024) and promote effective collaboration across different teams (Crisan et al., 2022; Gilbert et al., 2023; Heger et al., 2022; Mohammad, 2022; Pushkarna et al., 2022; Raji and Yang, 2020; Richards et al., 2020; Shen et al., 2022; Srinivasan et al., 2021).
@@
At a basic level, AI documentation could serve as an initial point of reference for downstream practitioners, offering an overview of the system that allows them to build foundational knowledge before engaging with developers for more in-depth discussions (Liao et al., 2023; Heger et al., 2022). Rather than acting as a comprehensive, self-contained repository of information, documentation artifacts could provide an initial layer of understanding, enabling downstream practitioners to identify relevant areas for further exploration and formulate more informed questions when collaborating with developers. This could foster more productive and targeted interactions between teams, enhancing their ability to work with or govern AI systems efficiently. Empirical research suggests that practitioners often approach documentation this way. For instance, UX professionals often prefer direct discussions with data scientists rather than relying solely on documentation artifacts (Liao et al., 2023). Similarly, many AI practitioners review existing documentation only briefly before seeking further clarification through meetings or discussions (Heger et al., 2022).
@@
In more robust forms, documentation can facilitate deeper deliberation among stakeholders about the system’s design and the documentation process itself (Pushkarna et al., 2022; Shen et al., 2022). For example, documentation has been shown to aid teams in collectively reflecting on socially constructed concepts, such as gender, and making informed decisions about how to annotate such attributes in datasets (Pushkarna et al., 2022). Also, the process of creating documentation can help stakeholders identify and discuss trade-offs between competing objectives in system design (Shen et al., 2022).
@@
4.2.3. Theory of change #3: Documentation could prompt practitioners to deliberate and act on the ethical impacts of AI
As with other tools that surface AI harms (Balayn et al., 2023), documentation could encourage practitioners to consider the ethical implications of their systems.
@@
One theory is that when downstream documentation consumers consult documentation on data, models, and systems, it can prompt consideration of potential harms. This consideration could lead practitioners to make more careful decisions about if and when to use certain systems or components. For example, one study found that practitioners with access to data documentation were more likely to identify ethical concerns unprompted. However, those without documentation often recognized similar concerns when prompted by the researcher. Despite this, practitioners both with and without access to documentation struggled to develop concrete action plans (Boyd, 2021). These findings indicate that while documentation can facilitate ethical deliberation, it is not always essential for ethical reflection and does not consistently lead to actionable solutions.
@@
Another theory is that documentation also might sensitize producers to the ethical impacts of their work because documentation frameworks often ask documentation producers to consider and describe the ethical dimensions of their systems. For example, the one framework prompts practitioners to think about privacy implications by inquiring about sensitive information in data (Gebru et al., 2021). Another asks about known biases, ethical issues, and safety concerns (Arnold et al., 2019), and several frameworks encourage reflection on how systems might negatively impact marginalized users or populations (Brajovic et al., 2023; McMillan-Major et al., 2021; Shen et al., 2022).
@@
4.2.4. Theory of change #4: Documentation could catalyze overall improvements in governance and development.
In applied settings, AI documentation rarely functions in isolation; it often relies on and reinforces complementary practices that could contribute to improved AI outcomes (Ahlawat et al., 2024). Our analysis revealed that several ways documentation could improve the rigor, efficiency, or reliability of development and governance practices overall.
@@
As one example, framework authors posit that documentation could improve the rigor with which practitioners develop AI systems (Holland et al., 2018). By requiring practitioners to justify their development choices, documentation may lead to more careful decision-making (Bender and Friedman, 2018; Chmielinski et al., 2024; Holland et al., 2018). Also, clear and comprehensive documentation supports reproducibility, aiding other practitioners in retraining and validating systems consistently (Adkins et al., 2022; Baracaldo et al., 2022; Hind et al., 2019). Since AI practitioners are often driven by a commitment to scientific rigor (Winecoff and Watkins, 2022) and product quality assurance (Ahlawat et al., 2024), documentation that emphasizes these aspects could potentially serve as a ”value lever (Shilton, 2013),” further encouraging engagement with the documentation process.
@@
Documentation could also facilitate proactive risk mitigation, which is often more efficient and effective than reactive approaches (Mitchell et al., 2019; Rakova et al., 2021). When issues arise in already-deployed systems, organizations typically need to fix the problem without disrupting service quality or availability. They might need to roll back to a previous version of the system that is less performant or apply quick patches that may not fully resolve the issue. Documentation that helps identify potential problems before deployment can lead to more comprehensive solutions than addressing issues post hoc once they are discovered in production systems (Holland et al., 2018; Ahlawat et al., 2024).
@@
Furthermore, documentation could preserve institutional knowledge. In any organization, particularly complex ones, consistently institutionalizing values and best practices can be challenging. Documentation could serve as a means of conveying technical information about systems, and the organization’s policies and values that influence development and use (Ahlawat et al., 2024). For instance, documentation can aid in onboarding new employees by providing information about the systems they’ll work on and communicating the organization’s approach to development and governance (Adkins et al., 2022; Chang and Custis, 2022; Heger et al., 2022).
@@
Lastly, documentation could facilitate both internal and external audits (Brajovic et al., 2023; Crisan et al., 2022; Hupont and Gomez, 2022; Miceli et al., 2021; Ahlawat et al., 2024). Audits may focus on the components of AI systems, the overall application, or the processes used in development (Mökander et al., 2023). Documentation could help organizations demonstrate that their claims—such as providing equitable system performance across different demographic groups—are backed by evidence. In these cases, documentation could guide auditors in selecting appropriate methods based on the auditor’s objectives and the available data (Ahlawat et al., 2024). When audits examine the robustness of an organization’s governance, documentation becomes even more critical since such information cannot be produced by analyzing the system itself (Clavell, 2024). Without comprehensive documentation, organizations may struggle to demonstrate the integrity and consistency of their processes.
4.3. Challenges
Both proposed frameworks and empirical studies delineated how documentation could support more robust governance of AI systems; however, empirical studies, in particular, illuminated challenges organizations and practitioners often face, which could prevent the theories of change from being fully realized. These challenges included limited incentives and resources for robust documentation, structural and organizational barriers to communication, interpersonal and organizational constraints to ethical action, and poor integration with existing workflows, practices, and organizational infrastructure.
4.3.1. Challenge #1: Poor incentivization and limited resources can undermine documentation quality.
Practitioners acknowledge that the benefits of documentation could outweigh the costs of creating and maintaining it (Heger et al., 2022); however, when organizations fail to provide clear incentives for high-quality documentation, practitioners are less likely to prioritize it. This lack of motivation can hinder the realization of documentation’s potential benefits.
Organizations typically do not prioritize documentation unless regulations require it for compliance or clients specifically request it (Chang and Custis, 2022; Miceli et al., 2021; Heger et al., 2022). Practitioners often view producing high-quality documentation as less relevant to their evaluations and promotions than development tasks that directly contribute to product outputs (Chang and Custis, 2022; Heger et al., 2022). This disconnect between the benefits of documentation and practitioner objectives is further exacerbated when the practitioners who benefit most from documentation (i.e., consumers) are not the same as those responsible for producing it (i.e., producers). In organizations where governance processes mandate documentation, a lack of familiarity with AI documentation concepts among developers could potentially magnify tensions between product development and governance teams (Ali et al., 2023; Ahlawat et al., 2024). Moreover, practitioners are often more motivated by compliance than by normative considerations, which could lead to blind spots when novel ethical issues arise outside existing policies (Heger et al., 2022).
Under time constraints and competing priorities, practitioners may rush the documentation process, sometimes cutting corners to save time (Chang and Custis, 2022; Heger et al., 2022). For instance, they may reuse content from existing documentation, even when it describes different systems, resulting in incomplete or inaccurate records for the current system (Pushkarna et al., 2022; Heger et al., 2022; Ahlawat et al., 2024). Practitioners also sometimes leave questions unanswered rather than seek out the necessary information (Heger et al., 2022) or prioritize recording information that is most valuable to them, potentially omitting critical details for others less familiar with the system (Hind et al., 2019) or leaving downstream practitioners with an incomplete end-to-end understanding of systems (Ahlawat et al., 2024). Furthermore, because practitioners often forget relevant information if they do not record it during development, their reconstruction of key system information after the fact can be time-consuming and error prone (Reid and Williams, 2023).
A potential consequence of rushed, low-quality documentation is that it may inadequately inform—or even misinform—downstream practitioners, potentially compromising effective development and risk management (Pepe et al., 2024; Liang et al., 2024; Geiger et al., 2020). For example, many publicly available model cards lack details about out-of-scope uses, limitations, or environmental impacts (Pepe et al., 2024; Liang et al., 2024). Less than 30% of model cards contain evaluation results, a critical piece of information for downstream practitioners (Liang et al., 2024). Also, some publicly available documentation contains incorrect license information, potentially leading practitioners to inadvertently violate license terms (Pepe et al., 2024).
Practitioners frequently advocate for automating as much of the documentation process as possible to maintain its relevance and reduce its time demands (Rostamzadeh et al., 2022; Heger et al., 2022; Pushkarna et al., 2022; Ahlawat et al., 2024). However, certain essential details—such as the rationale behind specific design or development decisions—are challenging to automate. This type of information is often crucial for evaluating potential harms and for enabling accountability (Yurrita et al., 2023; Sambasivan et al., 2021). Recognizing this, some practitioners acknowledged that parts of the documentation process must remain manual (Pushkarna et al., 2022). However, if these manually updated sections are not regularly synchronized with system updates, outdated information could itself become a source of harm (McMillan-Major et al., 2021). As a result, deprioritizing documentation could lead to a situation where the most critical elements for risk management are the most likely to become outdated.
4.3.2. Challenge #2: Structural and organizational barriers can complicate collaboration
Several practical barriers can limit how well documentation can act as a springboard for active communication and collaboration.
The modern AI supply chain, in which general-purpose models are adapted to a variety of downstream tasks (Vaswani et al., 2017), introduces additional structural complications for documentation. In some cases, different teams within the same organization may adapt a general-purpose model for specific tasks. In these instances, documentation might prompt the downstream team to directly reach out to the original development team to discuss the system’s capabilities and limitations or clarity details. However, in other cases, a team deploying a model may rely on a model developed by a different organization (Jones, 2023). In this case, it will typically be more difficult if not impossible for the downstream deployment team to directly communicate with the original development team. Any information that is missing or confusing in the documentation artifact cannot be easily clarified through collaboration.
Furthermore, when AI development implicates complex supply chain dynamics, actors further removed from the deployment context often struggle to account for the needs and considerations of external stakeholders (Balayn et al., 2024). This disconnect may manifest in risk documentation for upstream models that fails to accurately or comprehensively capture the real-world risks these systems pose in specific deployment contexts.
Challenges relevant to documentation of general-purpose models may be especially pronounced for less technical downstream stakeholders, who may struggle to understand the functionality and risks associated with these models (Crisan et al., 2022). More interactive forms of documentation, such as those that allow users to engage with the models directly or customize the level of detail presented, could help mitigate these issues by making the information more accessible and relevant (Crisan et al., 2022; Liao et al., 2023; Moore et al., 2023), especially when direct communication between teams is limited. However, further research is needed to determine the most effective ways to document general-purpose systems while avoiding overwhelming documentation consumers with excessive or irrelevant details.
Beyond the structural challenges of general-purpose models, a major organizational challenge is that documentation that is misaligned with the needs of its intended consumers can complicate rather than facilitate stakeholder communication. If documentation artifacts are overly technical or jargony, documentation consumers who are less technical or less familiar with the system may struggle to understand what details are most relevant, or misinterpret described practices (Crisan et al., 2022; Liao et al., 2023; Miceli et al., 2021; Shen et al., 2022); if documentation lacks relevant details, documentation consumers may overlook risks that become apparent only later in the development process, which may require teams to substantially backtrack. In either case, mismatch in comprehension can contribute to needless disruption to normal workflows.
4.3.3. Challenge #3: Interpersonal and organizational constraints can limit ethical decision making
In some cases, documentation can improve critical reflection (Boyd, 2021); however, several practical barriers may limit the extent to which documentation can improve ethical deliberation and decision making. First, the assumption that documentation artifacts and processes can promote ethical reasoning among practitioners hinges on the belief that practitioners are sufficiently aware of AI’s ethical risks to users, non-users, and society that they can recognize them when prompted. However, without proper training in responsible AI practices and exposure to groups potentially harmed by AI systems, practitioners may not connect documentation with ethical risks (Heger et al., 2022). Consequently, their consideration of harms may only partially encompass the scope of risks, especially those requiring non-technical mitigations. For example, practitioners often misunderstand how bias can manifest in their work, leading them to incorrectly state in their documentation that bias concerns are not relevant (Hind et al., 2019). While it may not be necessary for ethicists to be embedded in the work process for practitioners to make responsible choices, as some have suggested (McLennan et al., 2022), findings point to the need for more support and training for practitioners in how to identify, document, and effectively communicate potential risks (Heger et al., 2022; Madaio et al., 2024).
Second, practitioners and organizations may resist documenting ethical impacts they have identified. Publicly available documentation artifacts, such as Hugging Face model cards and GitHub repository README files, often show little consideration of ethical concerns (Pepe et al., 2024; Liang et al., 2024). Sometimes practitioners resist addressing ethical impacts because they feel unqualified to speculate on numerous potential use cases and their impacts (Crisan et al., 2022; Heger et al., 2022; Hind et al., 2019; Nunes et al., 2022; Ahlawat et al., 2024). Others worry that detailing ethical concerns might give downstream stakeholders a false sense of security (Crisan et al., 2022). This hesitation could also stem from a desire to avoid personal responsibility for their actions (Nunes et al., 2022). Alternatively, practitioners could plausibly have concerns that documenting potential ethical harms could create a legal or public relations risk if outside stakeholders gained access to documentation.
Third, even if practitioners recognize and record ethical concerns, they may have little influence over their organization’s development goals, deliverables, and timelines (Ahlawat et al., 2024). As a result, they might be unable to make the changes to data, models, or systems that they have identified as useful or necessary to address ethical risks through the documentation process (Ahlawat et al., 2024; Miceli et al., 2021). One ethnographic study found that business demands, not the beliefs of data subjects or practitioners, largely determined the organization’s documentation approaches (Miceli et al., 2021). For example, although practitioners who reflected on data labeling practices recognized that socially constructed identities are complex, the organization nevertheless chose to represent identity in a reductive way, such as by defining race according to discrete, mutually exclusive categories. The authors concluded that explicit and implicit power structures among internal organizational stakeholders significantly affect practitioners’ documentation practices and ability to shape outcomes. Documentation approaches must be responsive to these constraints. Otherwise, documentation about ethical considerations may not promote meaningful action or may be incomplete or misleading.
4.3.4. Challenge #4: Poor integration with technical and organizational infrastructure can undermine diffuse benefits
Practitioners studies emphasize that documentation is likely to be more effective when it is well integrated with practitioner workflows (Heger et al., 2022; Bhat et al., 2023). For example, when practitioners are asked to produce documentation using external tools, they are more likely to make errors and are less likely to make note of relevant ethical considerations than when documentation methods are well integrated with their existing tools (Bhat et al., 2023).